Fifteen carpenters can complete building a house in 45 days . How many extra men will be required if the job is to be completed in 25 days ?
Answer by ankor@dixie-net.com(22740)

Fifteen carpenters can complete building a house in 45 days.
Find the number of man-days required to complete the job
15*45 = 675
:
How many extra men will be required if the job is to be completed in 25 days?
let x = no.of extra men required
25(x+15) = 675
25x + 375 = 675
25x = 675 - 375
25x = 300
x = 300/25
x = 12 extra men
:
:
Check
25(15+12) = 675